Using Residue theorem evaluate ∫0^2π dθ/2+sinθ ?


1
Expert's answer
2022-01-14T09:41:38-0500

Make the substitution:"z = e^{ i\u03b8} , dz = ie^{i\u03b8}d\u03b8." Now the complex "z" describes the unit circle "C_1" in the positive sense as "\u03b8" varies from to "2\pi."

The integral becomes

"\displaystyle\int_{0}^{2\pi}\dfrac{d\theta}{2+\sin \theta}=\oint_{C_1}\dfrac{\dfrac{dz}{iz}}{2+\dfrac{z^2-1}{2iz}}"

"=\oint_{C_1}\dfrac{2dz}{z^2+4iz-1}"

The poles of the integrand are simple and occur when "z^2+4iz-1=0"


"z_{\pm}=\dfrac{-4i\pm
\sqrt{-16+4}}{2}=(-2\pm\sqrt{3})i"

"z_+z_-=(-2+\sqrt{3})i(-2-\sqrt{3})i=-1"

Therefore, "z_+" is a (simple) pole inside the unit circle.

Use "Res_{z=z_0}\dfrac{g(z)}{h(z)}=\dfrac{g(z_0)}{h'(z_0)}," where "z_0" is a simple zero of "h(z)."

Then

"Res(z_+)=\dfrac{2}{2z_{+}+4i}=\dfrac{1}{-2i+\sqrt{3}i+2i}=\dfrac{1}{\sqrt{3}i}"

Thus, by the residue theorem,


"\displaystyle\int_{0}^{2\pi}\dfrac{d\theta}{2+\sin \theta}=2\pi i(\dfrac{1}{\sqrt{3}i})=\dfrac{2\pi}{\sqrt{3}}"